i cannot figure out how to go about this problem
A jogger runs for 2.5 min(150 sec) at a constant speed of 5 m/s and then becomes tired, walks 50 m at a constant speed of 1.5 m/s. Then she runs another 250 m at a constant speed of 3.5 m/s. What is the joggers average speed over the whole motion.

Answer by KMST(5398)

The average speed is the total distance divided by the total time.
You are given the time, but not the distance for the first part of the trip. (You have to figure out the distance from the time and speed).
You are given the distances, but not the times for the rest of the trip. (You have to figure out the times from the distances and speeds).
Once you have the times and distances for each of the 3 parts of the trip, add to find the totals. Then divide to find the average speed.
Work that out on your own, as far as you can, then check my solution below and see if you agree. If not, one of us may have made a mistake. I often do.
(150 sec)x (5 m/s) = 750 m for the first portion,
(50 m)/(1.5 m/s) = (100/3) s = about 33.3 s for the middle part,
(250 m)/(3.5 m/s) = (500/7) s = about 71.4 s for the final portion of the trip.
The total time is
150 s + 33.3 s + 71.4 s = 254.7 s
The total distance is
750 m + 50 m + 250 m = 1050 m
(1050 m)/(254.7 s) = 4.1 m/s